What concentration of SO32- is in equilibrium with Ag2SO3(s) and 2.90 × 10-3 M Ag ? The Ksp of Ag2SO3 can be found here.
Answer #1
Concepts and reason

The solubility product for the reaction is equilibrium constant where the solid ionic compound dissociates into its ions in a solution. Solubility product is denoted as. Solubility product value relates to the saturated solution and indicates the precipitate level of the compound.

Fundamentals

Solubility product value of the compound depends on the concentrations of its ions in a solution.

Example: AB is any solid ionic compound.

AB(s) =A* (aq) +B* (aq)
K = A*B*]

Precipitation: If the solubility product value is lesser than the concentration of the ions present in the solution, the compound precipitates in the solution.

Given ionic solid is Ag,SO

Equilibrium equation of the ionic compound

Ag,SO,(s)=2Ag* (aq) +S0,(aq)

Now

Solubility product of the ionic compound

K. =[Ag*] [s0,?]

Given that

The concentration ion is 2.90x10 M

Solubility product of the given ionic solid is 1.5x10-14

Now

K. = [Ag*] [s0,]
1.5*10*4 =(2.90x10)[s0,]
[S0,]= _1.5*10-
(2.90x103)
=1.78x10M

Therefore, the concentration of is 1.78x10 PM

Ans:

The concentration of is 1.78x10 PM